Exact value of tan200(cot 10 – tan 10) is (all angles are in degree)

Grade:12

1 Answers

Samuel Garry
5 years ago
Tan 200 = tan (180 - 20) = tan 20
Tan 20( cot 10 - tan 10) => tan20((1-tan210)/tan 10)
2tan 10/1-tan10 (1-tan2 10)/tan 10) = 2.           Tan 2x= 2 tan a/ 1 - tana
 
ans: 2